The sum of the digits of a two-digit number is 11. If the digits are reversed, the new number is 45 more than the original number. Find the number.
Answer by stanbon(75887)

The sum of the digits of a two-digit number is 11. If the digits are reversed, the new number is 45 more than the original number. Find the number.
-------------
Let the two-digit number be 10t+u where t is the tens digit and u the units digit.
-------------------
EQUATION:
t+u = 11
10u+t = 10t+u + 45
-------------
Rearrange the equations:
t+u = 11
9t-9u = -45
-------------
Simplify:
t+u = 11
t-u = -5
---------
Add the equations to solve for "t":
2t = 6
t = 3
--------
Substitute to solve for "u":
3+u = 11
u = 8
--------------
The original number is 38
==============
Cheers,
Stan H.